• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name fansler

    Log entry time 21:51:06 on June14,2004

    Entry number 125764

    keyword=septa controls restored

    Septum magnets would not respond to user inputs.  Tried to reset IOC and PLC remotely with no success.  Performed hard reset on PLC, had no effect.  Reloaded Septum PLC software, and restored septum controls.  Magnets are ramping at this time.